OpenStudy (anonymous):

Can anyone set me a p=1 sphere bying using cylinderical corrdinates in integral.

OpenStudy (experimentx):

p = 1?? what is that??

OpenStudy (anonymous):

yeah same thing i wanna ask... what is p?

OpenStudy (anonymous):

rho=1

OpenStudy (experimentx):

what do you want to evaluate??

OpenStudy (anonymous):

sorry . the volume.

OpenStudy (experimentx):

looks like it's done here http://www.physicsforums.com/showthread.php?t=304505

OpenStudy (anonymous):

should it be this?

OpenStudy (anonymous):

\[\int\limits^{2\pi}_{0}\int\limits^{1}_{0}\int\limits^{\sqrt{1^2-r^2}}_{\sqrt{1^2-r^2}} rdz dr d\theta\]

OpenStudy (experimentx):

Hm ... i haven't tried yet, it will be good exercise for me

OpenStudy (anonymous):

u want to convert spherical coordinate into cylindrical coordinate?

OpenStudy (experimentx):

the upper and lower limit shouldn't be same

OpenStudy (anonymous):

yes @kr7210

OpenStudy (anonymous):

\[\int\limits\limits^{2\pi}_{0}\int\limits\limits^{1}_{0}\int\limits\limits^{\sqrt{1^2-r^2}}_{-\sqrt{1^2-r^2}} rdz dr d\theta\]lower s

OpenStudy (anonymous):

then first convert into cartesian coordinate thn go for cylindrical .....

OpenStudy (anonymous):

in this case r=rho

OpenStudy (anonymous):

r=rho=1

OpenStudy (anonymous):